the scores on the accuplacer test and High School GPAs are normally distributed. The Accuplacer test had a mean of 40 and a standard deviation of 10. High School GPAs had a mean of 2.5 and a standard deviation of 0.1. What high school GPA do you need to equal a score of 44 on the Accuplacer test? Give answer to two decimal places
Answer #1

Standardised score (z score) can be used to find an equivalent High School GPA score corresponding to the given Accuplacer test score.

z = (x - mean)/standard deviation

For Accuplacer test, mean = 40

Standard deviation = 10

When x = 44

z = (44 - 40)/10 = 0.4

This z score would be same for equivalent High School GPA

For High School GPA, mean = 2.5

Standard deviation = 0.1

0.4 = (x - 2.5)/0.1

x = 2.54